# DPER: Diffusion Prior Driven Neural Representation for Limited Angle and Sparse View CT Reconstruction

### v1

@@ 3.1 Overview
where Eq. (15b) denotes the prior subproblem, which estimates the noiseless image $\mathbf{x}_{0|t}$ via VE-SDE sampling with Tweedie’s formula. Eq. (15b) is the data-fidelity subproblem, and Eq. (15c) is the forward sampling operation by adding noise at timestep $t$.
For the data-fidelity subproblem as Eq. (15b), we employ INR to address it. The data-fidelity term $\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}$ ensures that the predicted $\mathbf{x}$ closely aligns with the measured data $\mathbf{y}$, thereby maintaining data fidelity. Additionally, the consistency term $\left\|\mathbf{x}-\mathbf{x}_{0|t}\right\|^{2}$ emphasizes the similarity of the predicted image $\mathbf{x}$ to the prior image $\mathbf{x}_{0|t}$, which serves to reduce artifacts that might arise from overfitting to incomplete measurements $\mathbf{y}$.
Fig. 1 shows the overview of the proposed DPER. In our approach, the distribution prior sub-problem is solved by the diffusion model as shown in Fig. 1(a). The image prior $\mathbf{x}_{T}$ is initially sampled from a Gaussian distribution $\mathcal{N}(0,1)$. Then, we generate a noisy prior image $\mathbf{x}_{t}$ at timestep $t$ by performing the reverse SDE process on a pre-trained score diffusion model. The generated noisy image $\mathbf{x}_{t}$ follows the pre-trained target distribution. Then we use Tweedie’s formula [43] to produce its denoised version $\mathbf{x}_{0\mid t}$, which serves as an initial solution to the distribution prior sub-problem (i.e., $\mathbf{z}_{k+1}=\mathbf{x}_{0\mid t}$ in Eq. (15a)). As demonstrated in Fig. 1(b), the initialized prior image $\mathbf{x}_{0\mid t}$ is subsequently transferred to the INR refinement phase to address the data fidelity sub-problem in Eq. (15b). In the prior embedding step (Fig. 1(b.1)), INR is initially trained to present $\mathbf{x}_{0\mid t}$. Then it is further refined towards the incomplete measurement $\mathbf{y}$ to generate a high fidelity and realistic solution as in Fig. 1(b.2). This refined image $\hat{\mathbf{x}}_{0\mid t}$ is the solution to the data fidelity sub-problem (i.e., $\mathbf{x}_{k+1}=\hat{\mathbf{x}}_{0\mid t}$ in Eq. (15b)). $\hat{\mathbf{x}}_{0\mid t}$ is then transferred back to the SDE process to further enhance the target distribution prior. By iteratively performing the reverse SDE sampling and INR refinement, the high-quality CT images can be reconstructed when the reverse SDE process is completed.
@@ 3.2 Generative Image Prior by Diffusion Model
We employ a score-based diffusion model of the ‘Variance Exploding (VE)’ SDE [44] form and utilize the reverse SDE sampling to address the distribution prior sub-problem in Eq. (15a). Specifically, the prior $\mathbf{x}_{T}$ is initialized with Gaussian distribution, and the denoising process based on score matching is executed sequentially with the timestep $T$ decay as follows:
@@
where $\sigma_{t}$ is the noise scale of $\mathbf{x}_{t}$. The expression is the same as the prior subproblem Eq. (15a) we derived in Sec. 3.1. Given the estimated score function provided by the pre-trained score model $\mathbf{s}_{\bm{\theta}^{*}}$, we employ Tweedie’s formula to get the noiseless prior image $\mathbf{x}_{0\mid t}$ from $\mathbf{x}_{t}$. $\mathbf{x}_{0\mid t}$ denotes the denoised result inferred from $\mathbf{x}_{t}$ for clarity and conciseness. The prior $\mathbf{x}_{0\mid t}$ is then employed to solve the data fidelity sub-problem with INR. The reverse SDE sampling corporating INR refinement to solve CT inverse problems can be rewritten specifically as:
$\mathbf{x}_{0\mid{t}}=\mathbf{x}_{t}+\sigma_{t}^{2}\mathbf{s}_{\bm{\theta}^{*}}\left(\mathbf{x}_{t},t\right)$ (18a)
$\hat{\mathbf{x}}_{0\mid t}=\mathbf{INR}(\mathbf{x}_{0\mid t},\mathbf{y})$ (18b)
${\mathbf{x}}_{t-1}=\hat{\mathbf{x}}_{0\mid t}+\sigma_{t-1}\bm{\epsilon},$ (18c)
where Eq. (18a) is the reverse diffusion sampling process with denoising by Tweedie’s formula, Eq. (18b) is the data fidelity sub-problem solving by INR (Sec. 3.3) and Eq. (18c) is the forward sampling operation by adding noise at timestep $t$.
@@ 3.3 Data Fidelity by Implicit Neural Representation
The data fidelity sub-problem in Eq. (15a) is addressed by the unsupervised deep learning architecture INR. Specifically, as shown in Fig. 1(b.2), we model the underlying CT image $\mathbf{x}$ as an implicit continuous function that maps image coordinates to intensities. Then CT imaging forward model is performed to simulate the physical acquisition process from the image $\mathbf{x}$ to its according projection measurements. By minimizing the errors on the measurement domain, we can optimize an MLP network $\mathcal{F}_{\mathbf{\Phi}}$ to approximate the implicit function of $\mathbf{x}$. The approximation of the implicit continuous function results in a highly accurate simulation of the CT imaging forward model, thereby ensuring effective data fidelity with the incomplete measurement $\mathbf{y}$.
@@ 3.3.1 Prior Embedding
In this stage, our goal is to incorporate the generative prior $\mathbf{x}_{0\mid t}$ from the diffusion model into the INR model. This strategy prevents INR from generating degradation reconstruction solution by avoiding overfitting to incomplete measurements $\mathbf{y}$. Technically, we represent the prior image $\mathbf{x}_{0\mid t}$ as a continuous function of image coordinate, which can be expressed as below:
@@
where $N$ denotes the number of sampling image coordinates at each iteration, $\mathbf{p}_{i}$ denotes the $i^{th}$ spatial coordinate. Once the optimization is completed, the prior image $\mathbf{x}_{0\mid t}(\mathbf{p})$ is embedded into the trainable weight $\mathbf{\Phi}_{\mathrm{prior}}$ of the MLP network. From the initialization perspective, we construct a feasible initialization that follows the target distribution for INR model using the generative prior provided by diffusion model. From an initialization standpoint, we establish a viable starting point for the INR model, utilizing the generative prior offered by the diffusion model.
@@ 3.3.2 Data Fidelity Refinement
$\mathbf{\Phi}_{\mathrm{refine}}$ $=\underset{\mathbf{\Phi}_{\mathrm{prior}}}{\mathrm{arg}\min}\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}|\hat{\mathbf{y}}(\mathbf{r})-\mathbf{y}(\mathbf{r})|,$ (21)
@@
where $\mathcal{R}$ denotes a set of sampling X-rays $\mathbf{r}$ at each iteration.
@@ 3.3.3 Image Reconstruction
11:         $\mathbf{\Phi}_{\mathrm{refine}}=\underset{\mathbf{\Phi}_{\mathrm{prior}}}{\mathrm{arg}\min}\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}|\hat{\mathbf{y}}(\mathbf{r})-\mathbf{y}(\mathbf{r})|$
@@ 3.5 Implementation Details
For the Diffusion Sampling stage (Fig. 1(a)), we set the total number of the sampling $T$ as 2000 and the refinement interval RI as 50. We adopt the ncsnpp [44] network as the score-based diffusion model architecture without any modification. For the INR refinement stage (Fig. 1(b)), we combine hash encoding [40] with five fully connected (FC) layers of 128 width to implement the MLP network $\mathcal{F}_{\Phi}$. The ReLU function is applied in each hidden layer. Hash encoding is a method which stores the trainable feature vectors in different multi-resolution hash tables to map the low-dimensional input to a higher dimensional space and thus enable network to efficiently learn higher frequency image feature. The hyper-parameters of the hash encoding are as follow: $L=10$, $T=2^{23}$, $F=8$, $N_{\text{min}}=2$, and $b=2$. During the prior embedding and data fidelity refinement, the numbers of training iterations are 50 and 250, the batch sizes are 1000 (i.e., $N=1000$ in Eq. (20)) and 128 (i.e., $|\mathcal{R}|=128$ in Eq. (21)), and the learning rates are 1e-3 and 1e-4. Adam optimizer with default hyper-parameters is leveraged. Note that all hyper-parameters are tuned on two samples of the AAPM dataset [48] and are then held constant across all other samples.
@@ 4.1.1 Dataset and Pre-Processing
We conduct experiments on two well-known public CT datasets, including AAPM 2016 low-dose CT grand challenge [48] and Lung Image Database Consortium (LIDC) image collection dataset [50]. The AAPM dataset DICOM Full Dose data was acquired at 120kV and 200mAs in the portal venous phase using a Siemens SOMATOM Flash scanner. For the acquired cone-beam projections, we approximated them to fan beam geometry using helix2fan, and then all volumes were reconstructed via standard filtered back projection (FBP). The LIDC database is not performed specifically for the purpose of the database so a heterogeneous range of scanner models (e.g., GE Medical Systems LightSpeed scanner models, Toshiba Aquilion scanners, Philips Brilliance scanner models) and technical parameters (e.g., 120-140kV tube peak potential energies, 40-627 mA tube current range, 0.6mm-5.0mm slice thickness) were represented. There are evident differences in the acquisition protocol and subjects between LIDC and AAPM datasets. Therefore, the evaluation based on the LIDC dataset can effectively reflect the generalization and robustness of the reconstruction methods in OOD settings. For the AAPM dataset, we extract 5376 2D slices from CT volumes of 9 subjects and randomly split them into 4240 training slices and 1136 validation slices. We then use 101 slices from an unseen volume as test data. For the LIDC dataset, we randomly choose 9 slices from various subjects as additional test data. Moreover, we retrospectively employ radon transform implemented by scikit-image library of Python (i.e., parallel X-ray beam) to generate the incomplete (SV and LA) projection data. Note that the training and validation images from only the AAPM dataset are used for training three supervised baselines (FBPConvNet, DCAR, and DDNet).
@@
We also conduct experimental validation on an in-house COVID-19 chest clinical CT dataset. All subjects were patients with severe COVID-19 who underwent serial chest CT scans during hospital stay. The examinations were performed with an Optima CT680 CT scanner (GE Medical Systems, USA;). The CT protocol was as follows: the tube voltage was 120 kVp, X-ray tube current was automatic (180$\sim$400 mA). The distance between the source and the detector was 94.9 cm, and the distance between the source and patient was 54.1 cm, respectively. The image matrix featured dimensions of 512$\times$512 pixels. The LACT and SVCT projections were extracted from the rebind fan-beam full-view projection data. To validate the robustness and performance of our method, we seamlessly applied the score function trained on the AAPM dataset to clinical data. It is worth noting that our score function was trained on 256$\times$256 images, so we generated a reconstruction of the same size and resized the reference image accordingly to compute the metrics.
@@ 4.1.2 Methods in comparison
DPER (Ours) 37.05$\pm$1.71 0.9587$\pm$0.0081 40.33$\pm$1.23 0.9722$\pm$0.0048 42.32$\pm$0.94 0.9846$\pm$0.0035
LIDC FBP[49] 10.20$\pm$1.09 0.1306$\pm$0.0207 15.30$\pm$0.90 0.2417$\pm$0.0303 18.52$\pm$0.83 0.3363$\pm$0.0375
FBPConvNet[11] 22.37$\pm$1.02 0.4623$\pm$0.0369 26.37$\pm$0.98 0.6081$\pm$0.3115 29.23$\pm$0.85 0.7417$\pm$0.0235
DDNet[13] 21.45$\pm$1.14 0.5209$\pm$0.0477 25.21$\pm$0.96 0.6238$\pm$0.0441 26.50$\pm$1.51 0.6809$\pm$0.0358
NeRP[27] 17.95$\pm$1.99 0.4894$\pm$0.0482 26.09$\pm$1.39 0.7327$\pm$0.0313 30.15$\pm$1.29 0.8299$\pm$0.0291
CoIL[25] 21.15$\pm$0.72 0.5209$\pm$0.0477 25.62$\pm$0.86 0.6270$\pm$0.0261 27.61$\pm$0.75 0.7251$\pm$0.0214
SCOPE[20] 22.28$\pm$0.69 0.6157$\pm$0.0498 26.87$\pm$0.89 0.7412$\pm$0.0436 29.56$\pm$1.14 0.8119$\pm$0.0424
MCG[34] 26.24$\pm$1.04 0.7679$\pm$0.0367 30.31$\pm$1.15 0.8623$\pm$0.0225 32.28$\pm$1.09 0.8897$\pm$0.0263
DiffusionMBIR[35] 26.75$\pm$0.57 0.8081$\pm$0.0306 31.04$\pm$1.02 0.8782$\pm$0.0286 33.48$\pm$1.19 0.9120$\pm$0.0221
DPER (Ours) 27.20$\pm$1.21 0.8318$\pm$0.0344 31.67$\pm$1.30 0.9005$\pm$0.0255 34.39$\pm$1.32 0.9321$\pm$0.0197

### v2

@@ 3.1 Overview
where Eq. (22a) denotes the prior subproblem, which estimates the noiseless image $\mathbf{x}_{0|t}$ via VE-SDE sampling with Tweedie’s formula. Eq. (22b) is the data-fidelity subproblem, and Eq. (22c) is the forward sampling operation by adding noise at timestep $t$. As mentioned earlier, the final convergence of the HQS-based penalty method to the optimal solution of the original problem presupposes that the penalty term $\mu$ gradually increases to $\infty$. In Eq. (22), the penalty term is $\left(\lambda/2\sigma_{t}^{2}\right)$, where $\lambda$ is a fixed scaling factor, and $\sigma_{t}^{2}$ decreases progressively according to an elaborate schedule and eventually approaching zero. This leads to a gradual increase of the penalty term to infinity, thus allowing our method to gradually converge to the optimal solution. Notably, our design enables the adjustment of penalty term weights $\mu$ to be intrinsically linked to the diffusion model’s sampling process. This approach eliminates the need for empirical adjustment of weights for different terms, thereby significantly enhancing the framework’s utility and robustness.
For the data-fidelity subproblem as Eq. (22b), we employ INR to address it. The data-fidelity term $\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}$ ensures that the predicted $\mathbf{x}$ closely aligns with the measured data $\mathbf{y}$, thereby maintaining data fidelity. Additionally, the consistency term $\left\|\mathbf{x}-\mathbf{x}_{0|t}\right\|^{2}$ emphasizes the similarity of the predicted image $\mathbf{x}$ to the prior image $\mathbf{x}_{0|t}$, which serves to reduce artifacts that might arise from overfitting to incomplete measurements $\mathbf{y}$. A more detailed discussion of this topic can be found in Sec. 3.3.
Fig. 1 shows the overview of the proposed DPER. In our approach, the distribution prior sub-problem is solved by the diffusion model as shown in Fig. 1(a). The image prior $\mathbf{x}_{T}$ is initially sampled from a Gaussian distribution $\mathcal{N}(0,1)$. Then, we generate a noisy prior image $\mathbf{x}_{t}$ at timestep $t$ by performing the reverse SDE process on a pre-trained score diffusion model. The generated noisy image $\mathbf{x}_{t}$ follows the pre-trained target distribution. Then we use Tweedie’s formula [60] to produce its denoised version $\mathbf{x}_{0\mid t}$, which serves as an initial solution to the distribution prior sub-problem in Eq. (22a). As demonstrated in Fig. 1(b), the initialized prior image $\mathbf{x}_{0\mid t}$ is subsequently transferred to the INR refinement phase to address the data fidelity sub-problem in Eq. (22b). In the prior embedding step (Fig. 1(b.1)), INR is initially trained to present $\mathbf{x}_{0\mid t}$. Then it is further refined towards the measurement $\mathbf{y}$ to generate a high fidelity and realistic solution as in Fig. 1(b.2). This refined image $\hat{\mathbf{x}}_{0\mid t}$ is the solution to the data fidelity sub-problem in Eq. (22b). $\hat{\mathbf{x}}_{0\mid t}$ is then transferred back to the SDE process to further enhance the target distribution prior. By iteratively performing the reverse SDE sampling and INR refinement, the high-quality CT images can be reconstructed when the reverse SDE process is completed.
@@ 3.2 Generative Image Prior by Diffusion Model
We employ a score-based diffusion model of the VE-SDE [62] form and utilize the reverse SDE sampling to address the distribution prior sub-problem in Eq. (22a). Specifically, the prior $\mathbf{x}_{T}$ is initialized with Gaussian distribution, and the denoising process based on score matching is executed sequentially with the timestep $T$ decay as follows:
@@
where $\sigma_{t}$ is the noise scale of $\mathbf{x}_{t}$. The expression is the same as the prior subproblem Eq. (22a) we derived in Sec. 3.1. Given the estimated score function provided by the pre-trained score model $\mathbf{s}_{\bm{\theta}^{*}}$, we employ Tweedie’s formula to get the noiseless prior image $\mathbf{x}_{0\mid t}$ from $\mathbf{x}_{t}$. $\mathbf{x}_{0\mid t}$ denotes the denoised result inferred from $\mathbf{x}_{t}$ for clarity and conciseness. The prior $\mathbf{x}_{0\mid t}$ is then employed to solve the data fidelity sub-problem with INR.
@@ 3.3 Data Fidelity by Implicit Neural Representation
The data fidelity sub-problem in Eq. (22b) is addressed by the unsupervised deep learning architecture INR. Specifically, as shown in Fig. 1(b.2), we model the underlying CT image $\mathbf{x}$ as an implicit continuous function that maps image coordinates to intensities. Then CT imaging forward model is performed to simulate the physical acquisition process from the image $\mathbf{x}$ to its according projection measurements. By minimizing the errors on the measurement domain, we can optimize an MLP network $\mathcal{F}_{\mathbf{\Phi}}$ to approximate the implicit function of $\mathbf{x}$. The approximation of the implicit continuous function between coordinates and the reconstructed image enables a highly accurate simulation of the CT imaging forward model, thereby ensuring effective data fidelity with the incomplete measurement $\mathbf{y}$.
@@ 3.3.1 Prior Embedding
In this stage, our goal is to incorporate the generative prior $\mathbf{x}_{0\mid t}$ from the diffusion model into the INR model. This strategy prevents INR from producing degradation reconstruction solution by avoiding overfitting to incomplete measurements $\mathbf{y}$. Technically, we represent the prior image $\mathbf{x}_{0\mid t}$ as a continuous function of image coordinate, which can be expressed as below:
@@
where $N$ denotes the number of sampling image coordinates at each iteration, $\mathbf{p}_{i}$ denotes the $i^{th}$ spatial coordinate. Once the optimization is completed, the prior image $\mathbf{x}_{0\mid t}(\mathbf{p})$ is embedded into the trainable weight $\mathbf{\Phi}_{\mathrm{prior}}$ of the MLP network.
From an initialization standpoint, we establish a viable starting point for the INR model, utilizing the generative prior offered by the diffusion model.
@@ 3.3.2 Data Fidelity Refinement
$\mathbf{\Phi}_{\mathrm{refine}}$ $=\underset{\mathbf{\Phi}_{\mathrm{prior}}}{\mathrm{arg}\min}\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}|\hat{\mathbf{y}}(\mathbf{r})-\mathbf{y}(\mathbf{r})|$ (27)
$+\frac{\lambda}{2\sigma_{t}^{2}}\left\|\mathcal{F}_{\mathbf{\Phi}}\left(\mathbf{p}\right)-\mathbf{x}_{0\mid t}\left(\mathbf{p}\right)\right\|^{2},$
@@
where $\mathcal{R}$ denotes a set of sampling X-rays $\mathbf{r}$ at each epoch.
@@ 3.3.3 Image Reconstruction
11:         $\mathbf{\Phi}_{\mathrm{refine}}=\underset{\mathbf{\Phi}_{\mathrm{prior}}}{\mathrm{arg}\min}\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}|\hat{\mathbf{y}}(\mathbf{r})-\mathbf{y}(\mathbf{r})|+\frac{\lambda}{2\sigma_{t}^{2}}\left\|\mathcal{F}_{\mathbf{\Phi}}\left(\mathbf{p}\right)-\mathbf{x}_{0\mid t}\left(\mathbf{p}\right)\right\|^{2}$
@@ 3.5 Implementation Details
For the Diffusion Sampling stage (Fig. 1(a)), we set the total number of the sampling $T$ as 2000 and the refinement interval RI as 50. We adopt the ncsnpp [62] network as the score-based diffusion model architecture without any modification. During the prior embedding and data fidelity refinement, the numbers of training iterations are 50 and 250, the batch sizes are 1000 (i.e., $N=1000$ in Eq. (26)) and 128 (i.e., $|\mathcal{R}|=128$ in Eq. (27)), and the learning rates are 1e-3 and 1e-4. Adam optimizer with default hyper-parameters is leveraged. Note that all hyper-parameters are tuned on two samples of the AAPM dataset [66] and are then held constant across all other samples.
@@ 4.1.1 Dataset and Pre-Processing
We conduct simulation experiments on two well-known public CT datasets, including AAPM 2016 low-dose CT grand challenge [66] and Lung Image Database Consortium (LIDC) image collection dataset [68]. The AAPM dataset DICOM Full Dose data was acquired at 120kV and 200mAs in the portal venous phase using a Siemens SOMATOM Flash scanner. The LIDC database is not performed specifically for the purpose of the database so a heterogeneous range of scanner models (e.g., GE Medical Systems LightSpeed scanner models, Toshiba Aquilion scanners, Philips Brilliance scanner models) and technical parameters (e.g., 120-140kV tube peak potential energies, 40-627 mA tube current range, 0.6mm-5.0mm slice thickness) were represented. There are evident differences in the acquisition protocol and subjects between LIDC and AAPM datasets. Therefore, the evaluation based on the LIDC dataset can effectively reflect the generalization and robustness of the reconstruction methods in OOD settings. For the AAPM dataset, we extract 5376 2D slices from CT volumes of 9 subjects and randomly split them into 4240 training slices and 1136 validation slices. We then use 101 slices from an unseen volume as test data. For the LIDC dataset, we randomly choose 30 slices from various subjects as additional test data. Moreover, we retrospectively employ radon transform implemented by scikit-image library of Python (i.e., parallel X-ray beam) to generate the incomplete (SV and LA) projection data. Note that the training and validation images from only the AAPM dataset are used for training three supervised baselines (FBPConvNet, DCAR, and DDNet).
Clinical Dataset with Lesion
@@
We also conduct experimental validation on an in-house COVID-19 chest clinical CT dataset. All subjects were patients with severe COVID-19 who underwent serial chest CT scans during hospital stay. The examinations were performed with an Optima CT680 CT scanner (GE Medical Systems, USA;). The CT protocol was as follows: the tube voltage was 120 kVp, X-ray tube current was automatic (180$\sim$400 mA). The distance between the source and the detector was 94.9 cm, and the distance between the source and patient was 54.1 cm, respectively. The image matrix featured dimensions of 512$\times$512 pixels. The LACT and SVCT projections were extracted from the rebind fan-beam full-view projection data. To validate the robustness and performance of our method, we seamlessly applied the score function trained on the AAPM dataset to clinical data.
Real Projection Dataset
In order to more closely match the real clinical scenario settings, we also conduct an experiment on the real projections data from the LDCT-and-Projection-data [^1] https://www.cancerimagingarchive.net/collection/ldct-and-projection-data/ [69]. It contains de-identified CT projection data and acquisition geometry collected in helical mode from either a GE Discovery CT750i, SOMATOM Definition AS+ or SOMATOM Definition Flash CT system. The tube potential is 80$\sim$120 kV, and the field-of-view (FOV) diameter covers an expansive 282$\sim$500 mm. Using the rebin technique described in [70], we obtain fan-beam projections with 2304 uniformly distributed scanning views over a complete 360-degree range. As with the COVID-19 dataset, the model trained on simulated data is seamlessly applied to the real projection data.
@@ 4.1.2 Methods in comparison
DPER (Ours) 38.16$\pm$0.82 0.9663$\pm$0.0054 41.06$\pm$0.85 0.9771$\pm$0.0043 42.32$\pm$0.94 0.9846$\pm$0.0035
LIDC FBP [67] 13.76$\pm$1.11 0.1609$\pm$0.0314 18.31$\pm$1.01 0.2854$\pm$0.0384 20.96$\pm$0.83 0.3465$\pm$0.0167
FBPConvNet [17] 25.64$\pm$0.80 0.6972$\pm$0.0244 30.91$\pm$1.11 0.8124$\pm$0.0211 34.05$\pm$1.09 0.8925$\pm$0.0148
DDNet[19] 24.62$\pm$0.87 0.7012$\pm$0.0239 28.81$\pm$0.92 0.8185$\pm$0.0178 29.44$\pm$0.50 0.8593$\pm$0.0124
NeRP[38] 20.92$\pm$1.34 0.5635$\pm$0.0569 27.33$\pm$1.36 0.7481$\pm$0.0421 31.74$\pm$2.13 0.8502$\pm$0.0474
CoIL[36] 25.63$\pm$1.83 0.8141$\pm$0.0344 29.79$\pm$1.18 0.8222$\pm$0.0381 32.19$\pm$1.08 0.8775$\pm$0.0262
SCOPE[31] 23.69$\pm$1.20 0.7052$\pm$0.0432 26.49$\pm$0.97 0.7682$\pm$0.0158 28.34$\pm$1.15 0.8662$\pm$0.0374
MCG[45] 30.33$\pm$1.01 0.8526$\pm$0.0187 33.81$\pm$0.99 0.8862$\pm$0.0211 35.48$\pm$1.14 0.9091$\pm$0.0144
DiffusionMBIR[46] 31.49$\pm$1.69 0.8730$\pm$0.0090 34.67$\pm$1.07 0.9069$\pm$0.0122 36.60$\pm$1.12 0.9396$\pm$0.0104
DPER (Ours) 32.52$\pm$1.11 0.9142$\pm$0.0133 36.10$\pm$1.18 0.9441$\pm$0.0047 37.79$\pm$1.02 0.9566$\pm$0.0137